A jet plane is coming in for a landing at 90.0 m/s. The wheels and the engines combined are able to decelerate the plane at a constant 1.85 m/s^2. How long of a runway does the plane require to come to a complete stop?
Answer #1

Initial speed, u = 90 m/s

acceleration, a = -1.85 m/s2 ( Negative sign is indicating that the plane is decelerating)

Final speed, v = 0 m/s

Distance covered, s = ?

We use the equation of motion

v2 - u2 = 2as

s = v2 - u2 / 2a = 02 - 902 / 2 (-1.85) m = 2189.2 m

Hence, the plane will run for 2189.2 m before stopping.